A line contains the point ( - 6, 7) and has a slope of . Write an equation of the line satisfying the given conditions. Write the answer in slope-intercept form.

and

A line contains the point ( - 3, - 4) and is perpendicular to a line with a slope of . Write an equation of the line satisfying the given conditions. Write the answer in slope-intercept form.

Answer by ikleyn(52781)

.

            Read your post and try to understand why it is nonsense.


Hey, one more such a post, and I will write a MAIL to your managers asking them to fire you from your position.



--------------
Comment from student: It can't be nonsense if its an actual question from school. its multiple choice but i refer to see how its broken down. These are the answers to choose from. If you can't help please do stop others from helping. A line contains the point ( - 3, - 4) and is perpendicular to a line with a slope of . Write an equation of the line satisfying the given conditions. Write the answer in slope-intercept form. A. y=-5/3x-4 B. y=-5/3x-9 C. y= 5/3x-9 D. y=-5/3x-3
--------------



MY response :


        Read your post and try to understand why it is nonsense.


                        Read your post and try to understand why it is nonsense.


                                        Read your post and try to understand why it is nonsense.